OpenStudy (anonymous):

If the angles A,B and C of a triangle are in arithmetic progression and if a,b and c denote the sides opposite to A, B and C respectively, then the value of the expression (a/c)Sin2C + (c/a)Sin2A is

OpenStudy (anonymous):

\[ \sin^2(C) \]or \[ \sin(2C) \]

OpenStudy (anonymous):

The latter

OpenStudy (anonymous):

Arithmetic progression just means:\[ B-A=C-B \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

This might help:\[ \frac{a\sin2C}{c}=\frac{a2\sin C\cos C}{c}=\frac{a2\sin A\cos C}{a} = 2\sin A\cos C \]

OpenStudy (anonymous):

Yes it has to do with the Sine and Cosine rule

OpenStudy (anonymous):

By symmetry, they both end up being the same thing, resulting in:\[ 4\sin A\cos C \]

OpenStudy (anonymous):

Wait that is wrong.

OpenStudy (anonymous):

\[ 2\sin A\cos C+2\sin C\cos A \]

OpenStudy (anonymous):

I found that in a IIT-jee paper

OpenStudy (anonymous):

i mean question

OpenStudy (anonymous):

Its the most difficult question in that paper they say

OpenStudy (anonymous):

We should probably use the \[ C=2B-A \]

OpenStudy (anonymous):

You continue to use angle formula

OpenStudy (anonymous):

One thing we could do is just say \[ A=30^\circ, B=60^\circ, C=90^\circ \]

OpenStudy (anonymous):

Wait i got to go. I think i will figure it out by tommorow and yes B=60 degrees